OpenStudy (anonymous):

Which of the following points are NOT on the graph of y = x + 2? A. (-1, 1) B. (1, 3) C. (3, 5) D. (5, 9)

OpenStudy (anonymous):

do you have a graphing calculator?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

what do you think the answer is

OpenStudy (anonymous):

c

OpenStudy (anonymous):

why do you think that?

OpenStudy (anonymous):

btw that's wrong

OpenStudy (anonymous):

because my mom help me

OpenStudy (anonymous):

y=x+2 c) (x,y)=(3,5) 5=3+2 5=5 so that means that point lies on the graph of y=x+2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

BUT you're question is asking which point is NOT on the graph

OpenStudy (anonymous):

ooooooooooo

OpenStudy (anonymous):

so what do you think the answr is now

OpenStudy (anonymous):

if you do what i did to the points given to you, and if one of the points shows that the left side does not equal the right side then you should have your answer

OpenStudy (anonymous):

a

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

(-1,1) = (x,y) y=x+2 1= -1+2 1 = 1

OpenStudy (anonymous):

now you're left with two, do the rest. that should be easy

OpenStudy (anonymous):

then i dont know ive been doing this all day

OpenStudy (anonymous):

yeah i know

OpenStudy (anonymous):

ok lol

OpenStudy (anonymous):

(5,9) is not on the graph. Because 9 is no t equal to 5+2

OpenStudy (anonymous):

thx